# Prekey Pogo: Investigating Security and Privacy Issues in WhatsApp’s Handshake Mechanism

### v1

@@ 1.1 Threat Model
We consider two different attack models: A PFS attack model where the attacker is assumed to have far-reaching capabilities, as well as a privacy and availability attack model where the sole requirement is having a WhatsApp account.
@@ 1.1.1 PFS Attack Model
Attacker Capabilities. In this attack, we assume that either a passive attacker has compromised WhatsApp’s data center internal communication or WhatsApp is forced to cooperate, s.t., an attacker is able to gain access to end-to-end encrypted communication of WhatsApp users. In other words, the attacker is able to strip the first layer of transport encryption — usually provided by TLS or the Noise protocol framework [24] — on top of the E2EE communication used in the Signal protocol suite. This is not an unrealistic scenario if a nation state actor, or compromised WhatsApp inter-server communication (e.g., by an internal employee), is considered. Moreover, minimizing the trust in the operator of the servers (i.e., WhatsApp in this case), is an explicit design goal of the Signal protocol family.
Prekey Depletion. Under this assumption, we describe the prekey depletion attack on a user Bob, in which the PFS of initial messages sent to him is violated by constantly depleting the ephemeral (one-time) prekeys, usually automatically deposited by Bob on the WhatsApp servers. We assume Bob to be a security-cautious user with a very strict deletion policy regarding messages (i.e., disappearing messages set to the lowest value; currently 24h). Therefore, a compromise of his long- and medium-term security keys usually would not lead to a compromise of message content, if PFS guarantees hold, as the plaintext of messages would no longer be available on his device.
@@ 1.1.2 Privacy and Availability Attack Model
Attacker Capabilities. For this attack model, the only requirement for the attacker is having a WhatsApp account.
@@ 2 Background
This section should provide a high level overview of the necessary protocol aspects of the Signal protocol. For more details we refer to Appendix 7.
@@
In this paper we focus on the desired (perfect) forwards secrecy (PFS) guarantees of the handshake protocol and the practical implications regarding privacy to ensure it. NIST defines perfect forward secrecy, or just forward secrecy for short, as follows:
@@
The Signal protocol uses three different types of Diffi-Hellman public keys to ensure forward secrecy right from the start: Long-term identity keys, medium-term (signed) prekeys and short-term (one-time) ephemeral prekeys (see Table 6 for an overview). In our scenario Alice is the initiator and wants to establish a secure connection with Bob (the responder). The Signal protocol, as also implemented by WhatsApp, uses prekey bundles deposited by every user at a central server to allow any initiator to negotiate a shared secret (via Diffi-Hellman Key exchange) even if the responder is currently not online/available using the X3DH protocol [20]  [^2] Since late 2023 Signal replaced X3DH by PQXDH [17]. On a high level, PQ3DH is comparable to X3DH with the difference that an additional key from a post quantum key encapsulation mechanism (KEM) is used..
Figure 1: High-level overview of the intended prekey bundle data deposit and retrieval process. Public keys end in $\dots pk$. Any key contains the creator, either Alice ($A$) or Bob ($B$), in superscript.
@@
The initial handshake works as depicted in Figure 1 and starting with Formula 1333For more details see Appendix 7.. First, the prekey bundle of the responder (in our case Bob) is fetched from the server by the initiator (in our case Alice). The information from the prekey bundle is verified by Alice through checking the signature on the signed prekey using the (long-term) identity pubic key of Bob. As within other works [8], it is assumed that Alice has already verified out-of-band that the long-term identity public key indeed belongs to Bob.
Then the public keys from Bob’s prekey bundle are used to compute shared keys for the ratcheting and message encryption and authentication. Here now we have to distinguish between the case where an ephemeral (one-time) prekey $eprepk^{B}$ of Bob is available or not. If no ephemeral prekey is available, the DH invocation $dh_{4}$ in Formula 20 is omitted.
@@
$\vdots$ (6)
At the end, a new message key $mk$ is derived using a key derivation function (KDF). This message key is then used to encrypt and authenticate ($AE$) the first chat message from Alice to Bob, as well as to authenticate some associated data $AD$ consisting of the ephemeral public keys generated previously by Alice.
@@
Once Bob receives this initial message, he can compute the same shared keys using his identity key ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}ik^{B}}$ and his prekey ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}prek^{B}}$, as well as the public keys of Alice consisting of her identity key $ipk^{A}$, her ephemeral handshake public keys. Since the later are transmitted in the associated data $AD$, they are authenticated, but not encrypted.
@@ 3.1 Relevant Endpoints and Message Structs
44footnotetext: Baileys: github.com/WhiskeySockets/Baileys
@@ 4 Results and Exploitation
Table 3: Success rate among different devices, i.e., how likely it is that a prekey bundle fetched for a new session will not contain a one-time prekey if the target is currenlty online and an attacker is actively depleting its one-time prekeys.
@@ 4.1 Perfect Forward Secrecy Degradation
For this section, we consider the PFS attack model described in Section 1.1.1, so the goal G1 of the attack is it to degrade PFS for new sessions initiated by other users (such as Alice). Therefore, Eve wants to deplete all one-time prekeys of Bob.
For our first attack we consider the simplest case, where the targeted device is currently offline and thus cannot refill depleted one-time prekeys at the moment. This attack is depicted in Figure 2. Our tests showed, since WhatsApp does not enforce any rate limiting, constantly querying prekey bundles for any user is possible, thereby eventually depleting all available one-time prekeys. Although one-time prekeys are crucial for ensuring PFS, messages can also be transmitted when only identity and signed prekey are available to generate a shared secret (i.e., when no one time prekeys are available on the server). In our tests, this PFS degradation is not indicated to the users UI (e.g. by a security notification in the WhatsApp client initiating the session). Therefore, the PFS degradation attack can be executed without the affected users noticing in their UI.
@@
If the respective device is online, it receives a notification from the server as soon as there are less than 11 one-time prekeys left. Depending on the device type and its current state, it reacts upon this notification and uploads 812 new one-time prekeys to the server. For measurements on how fast depletion of one-time prekeys is possible in case the targeted device is online, as well as the detailed behavior and availability implications, see Section 4.4 and 4.5. Table 3 provides an overview of the measured success probabilities of the one-time prekey depletion attack in case the targeted device is online.
@@ 5 Related Work
Security and Privacy at Instant Messaging. Schrittwieser et al. were the first to investigate security and privacy issues in mobile instant messaging and VoIP applications, uncovering vulnerabilities such as account hijacking and number spoofing [27, 22]. Beyond Over-the-Top (OTT) applications, similar security vulnerabilities have been identified in VoIP-based messaging solutions like VoLTE, VoWiFi and RCS [28, 15, 14, 31]. In many cases, these vulnerabilities remained undiscovered for years due to proprietary clients and the lack of tooling for security experiments. We adopt a similar security testing approach to find vulnerabilities within WhatsApp, leveraging open-source tools to emulate a real client sending protocol queries to the server. In contrast, Hagen et al. [16] demonstrated that large-scale account enumeration is possible in major messaging applications (e.g., WhatsApp, Signal) simply by automating interactions with the regular user interface.
@@ 6.1 Ethical Considerations
For our measurements and during experimentation we only targeted WhatsApp accounts under our direct control. Additionally, we tried to adhere to WhatsApp’s protocol through the use of community-proven open source projects (some of them being used in widely deployed production systems  [^8] https://github.com/element-hq/mautrix-whatsapp). In our depletion experiments, we issued a substantially higher volume of prekey queries compared to typical client implementations. However, this increased traffic is unlikely to pose a significant risk to WhatsApp’s infrastructure, which is built to serve more than three billion users. Moreover, we limited our offensive depletion to at most two concurrent client sessions. Lastly, none of our testing accounts were blocked throughout the study, hinting that we did not cause any significant harm and most likely were not even noticed by the platform operator. Lastly, to accurately assess the feasibility of the proposed attack, it was essential to conduct experiments against the actual WhatsApp infrastructure. Given the minimal risk of adverse effects on other users or the service itself, as argued above, we considered this a reasonable approach. Finally, all our findings will be responsibly disclosed to Meta, to give them enough time to fix the described vulnerabilities prior to publication.
@@ 6.3 Countermeasures
Reduce Signed Prekey Renewal Interval. The lifetime of a singed prekey in WhatsApp is higher ($\approx$ month), than the lifetime of a signed prekey in Signal (two days). Reducing the lifetime of signed prekeys would also reduce the impact regarding PFS through missing one-time prekeys.
@@ 7 Conclusion
Tested Phone Models and OS Versions
@@
Signal Protocol in More Detail
@@
The initial handshake works as depicted in figure 5 in the appendix and here starting with formula 16. First the prekey bundle of the responder (in our case Bob) is fetched from the server by the initiator (in our case Alice). The information from the prekey bundle is verified by Alice through checking the signature on the signed prekey using the (long-term) identity pubic key of Bob. As within other works [8], it is assumed that Alice has already verified out-of-band that the long-term identity public key indeed belongs to Bob.
Then the public keys from Bob’s prekey bundle are used to compute shared keys for the ratcheting and message encryption and authentication. Here now we have to distinguish between the case where an ephemeral (one-time) prekey $eprepk^{B}$ of Bob is available or not. If no ephemeral prekey is available, the DH invocation $dh_{4}$ in formula 20 is omitted.
@@
If no ephemeral prekeys have been fetched by Alice, this initial message sent by Alice has no forward secrecy if observed by an attacker. Therefore, an attacker who is able to compromise Bobs medium-term and long-term secret keys $prek^{B}$ and $ik^{B}$ later on, can recompute the same message key $mk_{0,0}^{i\to r}$, which highlights that there is no forward secrecy for this message. If Alice sends multiple messages before receiving any response from Bob, all these messages are affected as well, as the keys for these messages come from the symmetric ratchet. This is illustrated by the following example starting with formula 34, which depicts encrypting a second message from Alice to Bob. Here, $y$ is 00 at the beginning and later set to $y=1$ for the second message and so forth:

### v2

@@ 1.1 Threat Model
We consider two different attack models: A PFS attack model where the attacker is assumed to have far-reaching capabilities, as well as a privacy and availability attack model where the sole requirement for the attacker is having a WhatsApp account and the phone number of the target.
@@ 1.1.1 PFS Attack Model
Attacker Capabilities. In this attack, we assume a passive attacker that has access to WhatsApp’s data center internal network communication, s.t., they are able to gain access to end-to-end encrypted messages of WhatsApp users. In other words, the attacker is able to strip the first layer of transport encryption – usually provided by TLS or the Noise protocol framework between the client and the server [26] – on top of the E2EE communication used in the Signal protocol suite. This is not an unrealistic scenario if a nation state actor, or compromised WhatsApp inter-server communication (e.g., by an internal employee), is considered. Moreover, minimizing the trust in the operator of the servers (i.e., WhatsApp in this case), is an explicit design goal of the Signal protocol family [8]. It is therefore reasonable to assume that an attacker could, under certain circumstances, gain access to end-to-end encrypted (E2EE) messages.
Moreover, we require the attacker to know the phone number of the target user, as well as a WhatsApp account.
Prekey Depletion. Under these assumptions, we describe the prekey depletion attack on a user Bob, in which the PFS of initial messages sent to him is violated by constantly depleting the ephemeral (one-time) prekeys, usually automatically deposited by Bob on the WhatsApp servers. We assume Bob to be a security-cautious user with a very strict deletion policy regarding messages (i.e., disappearing messages set to the lowest value; currently 24h). Therefore, a compromise of his long- and medium-term security keys usually would not lead to a compromise of message content, if PFS guarantees hold, as the plaintext of messages would no longer be available on his device.
@@
Note that this paper focuses exclusively on point-to-point communication and does not consider group messaging scenarios.
@@ 1.1.2 Privacy and Availability Attack Model
Although the objectives of the two attack models differ, both exploit the same underlying prekey mechanisms and functionalities. As a result, the attacks are closely related and interconnected, underscoring the inherent privacy and security challenges associated with this mechanism.
@@
Attacker Capabilities. For this attack model, the only requirements for the attacker are knowing the telephone number of the target and having a WhatsApp account.
@@ 2 Background
This section should provide a high level overview of the necessary protocol aspects of the Signal protocol. For more details we refer to Appendix B.
@@
In this paper we focus on the desired (perfect) forward secrecy (PFS) guarantees of the handshake protocol and the practical implications regarding privacy to ensure it. NIST defines perfect forward secrecy, or just forward secrecy for short, as follows:
@@
The Signal protocol uses three different types of Diffi-Hellman public keys to ensure forward secrecy right from the start: Long-term identity keys, medium-term (signed) prekeys and short-term (one-time) ephemeral prekeys (see Table 1 for an overview). In our scenario Alice is the initiator and wants to establish a secure connection with Bob (the responder). The Signal protocol, as also implemented by WhatsApp, uses prekey bundles deposited by every user at a central server to allow any initiator to negotiate a shared secret (via Diffi-Hellman Key exchange) even if the responder is currently not online/available using the X3DH protocol [22]  [^2] Since late 2023 Signal replaced X3DH by PQXDH [19]. On a high level, PQXDH is comparable to X3DH with the difference that an additional key from a post quantum key encapsulation mechanism (KEM) is used..
Figure 1: High-level overview of the intended prekey bundle data deposit and retrieval process, illustrating the ideal case in which everything functions as expected (i.e., prekey bundle of Bob contains also a one-time prekey).
@@
Table 1: Main cryptographic keys of the signal protocol relevant for our attacks. Public keys in asymmetric schemes always end in $pk$. The naming convention of the keying material is according to Cohn-Gordon et al. [9]. The ephemeral prekeys (one-time prekeys), which are considered optional in the prekey bundle, are depicted in red. This is the key type drained in our attack. The secret keys of Bob which an attacker has to compromise to benefit from the violation of forward secrecy, i.e., if no ephemeral prekeys can be used, are depicted in orange.
The initial handshake works as depicted in Figure 1 and starting with Formula 1. For more details we refer the reader to the Appendix B. First, the prekey bundle of the responder (in our case Bob) is fetched from the server by the initiator (in our case Alice). The information from the prekey bundle is verified by Alice through checking the signature on the signed prekey using the (long-term) identity pubic key of Bob. As within other works [9], it is assumed that Alice has already verified out-of-band that the long-term identity public key indeed belongs to Bob.
Then the public keys from Bob’s prekey bundle are used to compute shared keys for the ratcheting and message encryption and authentication. Here now we have to distinguish between the case where an ephemeral (one-time) prekey $eprepk^{B}$ of Bob is available or not. If no ephemeral prekey is available, the DH invocation $dh_{4}$ in Formula 5 is omitted. Since this is the key type drained in our attack it is colored in red.
@@
$\vdots$
At the end, a new message key $mk$ is derived using a key derivation function (KDF). This message key is then used to encrypt and authenticate ($AE$) the first chat message from Alice to Bob, as well as to authenticate some associated data $AD$ consisting of the ephemeral public keys generated previously by Alice. This is shown in Formula 6 and 7.
@@
Once Bob receives this initial message, he can compute the same shared keys using his identity key ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}ik^{B}}$ and his prekey ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}prek^{B}}$, as well as the public keys of Alice consisting of her identity key $ipk^{A}$ (Formular 8) and her ephemeral handshake public keys (Formular 9 and 10). Since the later are transmitted in the associated data $AD$, they are authenticated, but not encrypted. Tho benefit from the PFS violation (due to missing one-time prekys) and successfully decrypt recorded E2EE messages, the orange keys have to compromised by the attacker later on.
@@ 3.1 Relevant Endpoints and Message Structs
33footnotetext: Baileys: github.com/WhiskeySockets/Baileys
@@ 4 Results and Exploitation
Table 3: Success rate among different devices, i.e., how likely it is that a prekey bundle fetched for a new session will not contain a one-time prekey even if the target is currenlty online and powered on and an attacker is actively depleting its one-time prekeys.
@@ 4.1 Perfect Forward Secrecy Degradation
For this section, we consider the PFS attack model described in Section 1.1.1, so the goal G1 of the attack is it to degrade PFS for new sessions initiated with Bob by other users (such as Alice). Therefore, Eve wants to deplete all one-time prekeys of Bob.
For our first attack we consider the simplest case, where the targeted device is currently offline and thus cannot refill depleted one-time prekeys at the moment. This attack is depicted in Figure 2. Our tests showed, since WhatsApp does not enforce any rate limiting, constantly querying prekey bundles for any user is possible, thereby eventually depleting all available one-time prekeys. Although one-time prekeys are crucial for ensuring PFS, messages can also be transmitted when only identity and signed prekey are available to generate a shared secret (i.e., when no one-time prekeys are available on the server). In our tests, this PFS degradation is not indicated to the users UI (e.g. by a security notification in the WhatsApp client initiating the session). Therefore, the PFS degradation attack can be executed without the affected users noticing in their UI. Note that even if Bob’s device comes online after Alice has initiated a new communication session without a one-time prekey, messages sent by Alice will not achieve PFS until Bob’s device responds for the first time within that session.
@@
Now let’s consider the case where the targeted device of Bob is constantly online. In this case, it receives a notification from the server as soon as there are less than 11 one-time prekeys left. Depending on the device type and its current power saving state, it reacts upon this notification and uploads 812 new one-time prekeys to the server. For measurements on how fast depletion of one-time prekeys is possible in case the targeted device is online, as well as the detailed behavior and availability implications, see Section 4.4 and 4.5. Table 3 provides an overview of the measured success probabilities of the one-time prekey depletion attack in case the targeted device is online. In other words, it indicates the likelihood that another party (e.g., Alice) will initiate a new communication session with Bob without using a one-time prekey while our depletion attack is in progress.
@@ 4.7 Peripheral Observations
Blocked Contacts. Because the attacker interacts only with the server and never sends direct packets to the victim, the victim has no means of identifying the source of the attack. Nevertheless, we evaluated whether blocking an account in WhatsApp has any impact by testing a scenario in which the victim had already blocked the attacker prior to the attack. The results show that blocking has no effect, thus the attack remains fully feasible.
@@ 5 Related Work
Security and Privacy at Instant Messaging. Schrittwieser et al. were the first to investigate security and privacy issues in mobile instant messaging and VoIP applications, uncovering vulnerabilities such as account hijacking and number spoofing [29, 24]. Beyond Over-the-Top (OTT) applications, similar security vulnerabilities have been identified in VoIP-based messaging solutions like VoLTE, VoWiFi and RCS [30, 17, 16, 33]. In many cases, these vulnerabilities remained undiscovered for years due to proprietary clients and the lack of tooling for security experiments. We adopt a similar security testing approach to find vulnerabilities within WhatsApp, leveraging open-source tools to emulate a real client sending protocol queries to the server. In contrast, Hagen et al. [18] demonstrated that large-scale account enumeration is possible in major messaging applications (e.g., WhatsApp, Signal) simply by automating interactions with the regular user interface. Beyond that, efficient account enumeration can also be achieved by using open-source clients to directly extract data from internal user APIs [14].
@@ 6.1 Ethical Considerations
For our measurements and during experimentation we only targeted WhatsApp accounts under our direct control. Additionally, we tried to adhere to WhatsApp’s protocol through the use of community-proven open source projects (some of them being used in widely deployed production systems  [^7] https://github.com/element-hq/mautrix-whatsapp). In our depletion experiments, we issued a substantially higher volume of prekey queries compared to typical client implementations. However, this increased traffic is unlikely to pose a significant risk to WhatsApp’s infrastructure, which is built to serve more than three billion users. Moreover, we limited our offensive depletion to at most two concurrent client sessions. Lastly, none of our testing accounts were blocked throughout the study, hinting that we did not cause any significant harm and most likely were not even noticed by the platform operator. Lastly, to accurately assess the feasibility of the proposed attack, it was essential to conduct experiments against the actual WhatsApp infrastructure. Given the minimal risk of adverse effects on other users or the service itself, as argued above, we considered this a reasonable approach. Finally, all our findings have been responsibly disclosed to Meta.
@@ 6.3 Countermeasures
Reduce Signed Prekey Renewal Interval. The lifetime of a signed prekey in WhatsApp is higher ($\approx$ month), than the lifetime of a signed prekey in Signal (two days). Reducing the lifetime of signed prekeys would also reduce the impact regarding PFS through missing one-time prekeys.
@@ 7 Conclusion
Artifact Evaluation
The artifact accompanying this paper is publicly available at https://github.com/sbaresearch/prekey-pogo/tree/woot25ae. It includes a modified WhatsApp client based on whatsmeow, which enables the retrieval of a user’s WhatsApp devices and their associated prekey material. As part of the WOOT 2025 artifact evaluation process, we applied for and were awarded both the Artifacts Available and Artifacts Functional badges.
Acknowledgments
This material is based upon work partially supported by (1) the Christian-Doppler-Laboratory for Security and Quality Improvement in the Production System Lifecycle; The financial support by the Austrian Federal Ministry for Digital and Economic Affairs, the Nation Foundation for Research, Technology and Development and University of Vienna, Faculty of Computer Science, Security & Privacy Group is gratefully acknowledged; (2) the FFG Bridge project 46322124 SecKey; (3) SBA Research (SBA-K1 NGC), a COMET Center within the COMET – Competence Centers for Excellent Technologies Programme, funded by BMIMI, BMWET, and the federal state of Vienna. The COMET Programme is managed by FFG.
We further would like to thank Markus Maier for his outstanding support in providing and maintaining the IT infrastructure, which was essential for the success of this work. Moreover, we would also like to thank our anonymous reviewers for their valuable feedback and suggestions.
@@ Acknowledgments
A Tested Phone Models and OS Versions
@@
B Signal Protocol in More Detail
@@
The initial handshake works as depicted in figure 5 in the appendix and here starting with formula 13. First the prekey bundle of the responder (in our case Bob) is fetched from the server by the initiator (in our case Alice). The information from the prekey bundle is verified by Alice through checking the signature on the signed prekey using the (long-term) identity pubic key of Bob. As within other works [9], it is assumed that Alice has already verified out-of-band that the long-term identity public key indeed belongs to Bob.
Then the public keys from Bob’s prekey bundle are used to compute shared keys for the ratcheting and message encryption and authentication. Here now we have to distinguish between the case where an ephemeral (one-time) prekey $eprepk^{B}$ of Bob is available or not. If no ephemeral prekey is available, the DH invocation $dh_{4}$ in formula 17 is omitted.
@@
If no ephemeral prekeys have been fetched by Alice, this initial message sent by Alice has no forward secrecy if observed by an attacker. Therefore, an attacker who is able to compromise Bobs medium-term and long-term secret keys $prek^{B}$ and $ik^{B}$ later on, can recompute the same message key $mk_{0,0}^{i\to r}$, which highlights that there is no forward secrecy for this message. If Alice sends multiple messages before receiving any response from Bob, all these messages are affected as well, as the keys for these messages come from the symmetric ratchet. This is illustrated by the following example starting with formula 31, which depicts encrypting a second message from Alice to Bob. Here, $y$ is 00 at the beginning and later set to $y=1$ for the second message and so forth: